OpenStudy (anonymous):

3a-3b=-15 -3a-3b=-3 solve using elimination method

OpenStudy (anonymous):

I am thinking like terms so 3 and -3 cancel eachother?

OpenStudy (diyadiya):

yes :)

OpenStudy (anonymous):

add the two equation to eliminate a

OpenStudy (diyadiya):

3a-3b+(-3a-3b)=-15-3 -3b-3b=-18 -6b=-18 b=3

OpenStudy (diyadiya):

3a-3b=-15 3a-3(3)=-15 3a=-15+9 3a=-6 a=-2

OpenStudy (anonymous):

Thanks, I did it a little different but somehow I got the right answers

OpenStudy (anonymous):

Thanks, I did it a little different but somehow I got the right answers

OpenStudy (anonymous):

Thanks, I did it a little different but somehow I got the right answers